For the given pair of functions, find the requested function.
f(x)=x^2+3x-2, g(x)=-9^2+9x-7; (f+g)(x)
I just don't understand this problem at all.
Thank you w_ann

Answer by stanbon(48535)

For the given pair of functions, find the requested function.
f(x)=x^2+3x-2, g(x)=-9x^2+9x-7;
-----------------------------------
(f+g)(x) = (x) + g(x)
= -8x^2 +12x -9
====================
Cheers,
Stan H.